if 1/2 is subtracted from 5 times the reciprocal of a number x, the result is 2, what equqation could be used to find x?
can you explain me please step by step thank you.

Answer by stanbon(75887)

If 1/2 is subtracted from 5 times the reciprocal of a number x, the result is 2, what equation could be used to find x?
---
Equation:
5(1/x) - (1/2) = 2
-----
Multiply thru by 2x to get:
10 - 5x = 4x
10 = 9x
x = (10/9)
---
x = 1 1/9
===============
Cheers,
Stan H.
====================